OpenStudy (carolinar7):

How to graph 3x^2+12x+16

OpenStudy (carolinar7):

Nnesha

OpenStudy (carolinar7):

@alexistheking777

OpenStudy (anonymous):

okay well theirs a couple of ways of doing this

OpenStudy (carolinar7):

ok

OpenStudy (anonymous):

one is just plugging in x values and plotting them then connecting the dots and presto

OpenStudy (carolinar7):

OK

OpenStudy (carolinar7):

I get It now

OpenStudy (carolinar7):

another question though

OpenStudy (anonymous):

ask away

OpenStudy (carolinar7):

how do I find the minimum value of that without graphing

OpenStudy (anonymous):

okay well you could find the derivative and then set that equal to zero

OpenStudy (carolinar7):

How do I do that

OpenStudy (anonymous):

are you taking calculus or algebra?

OpenStudy (carolinar7):

Algebra 2

OpenStudy (anonymous):

okay so what you want to find is the vertex and that's different to find the vertex its easier if you set your equation to vertex form

OpenStudy (carolinar7):

And how does the vertex for look like appliying the values that I gave u

OpenStudy (anonymous):

To convert a quadratic from y = ax2 + bx + c form to vertex form, y = a(x - h)2+ k, you use the process of completing the square

OpenStudy (carolinar7):

I dont know completing the square